# A two-phase model of galaxy formation: III. The formation of globular clusters

### v1

@@ 4.6 Spatial distribution
Figure 14: Projected spatial distribution of GCs within a MW-size galaxy with $M_{\rm v}=10^{12}\,h^{-1}{\rm M_{\odot}}$ at $z=0$. All galactic GCs predicted by our model and resampled with the BSampling method (see §4.1) are shown. Each dot represents a survived GC, with bold and thin edges indicating Pop-I and Pop-II GCs (see §3.5), respectively, The size of each circle is proportional to $R_{\rm cls}$, the half-mass radius of the GC (see §3.4). Metallicity of each GC is color-coded according to the color bar on the top. The big circle indicates two times the exponential truncation radius of the in-situ, slow-phase GC system of this galaxy. See §4.6 for a detailed discussion of this figure.
Figure 15: Number density profile of galactic GCs. Three panels show the results for central galaxies with different halo masses at $z=0$. Both in-situ and ex-situ components are included. In each panel, black, red and blue solid curves are for all, Pop-I and Pop-II GCs, respectively, all including only survived GCs. Black dashed curve includes all GCs at formation (see §4.1 for the selection). Each curve shows the median of GC systems, with error bars or shaded areas indicating the 1-$\sigma$ range. In the center panel, green curves are obtained by Rodriguez et al. (2023) using the model of GC formation developed by Grudić et al. (2023) and a star-by-star model of cluster evolution, applied to a MW-size galaxy in the FIRE-2 simulation. The solid and dashed curves include their GCs at birth and survived at $z=0$, respectively. Orange curve is obtained by their compilation of MW GCs based on the catalog of Harris (1996, 2010). See §4.6 for a detailed discussion of this figure.
@@
where GCs belonging to the same component share a common parameter $r_{\rm t}$ that describes the characteristic size of the distribution of the GC component. The isothermal form, $1/r^{2}$, is motivated by the expectation that the fragmentation of an SGC is associated with effective atomic cooling that keeps the temperature to be $10^{4}\,{\rm K}$ at the formation of dense sub-clouds (§3.1). The exponential truncation ensures that the distribution of galactic GCs has a finite extent. Once $r$ is sampled for a GC, the angular location is randomly drawn from a uniform distribution on a unit sphere.
@@
Fig. 15 shows the number density profiles of GCs in central galaxies with different halo masses at $z=0$. For the profile of all survived GCs (black solid curve), the amplitude increases rapidly with halo mass, as a result of the tight, linear relation between $M_{*}^{\rm(GCs)}$ and $M_{\rm v}$ described in §4.4. In contrast, the shape of the profile is independent of halo mass, which is a consequence of our model to assign positions to GCs (see Eq. 74). Such a universal profile has been found in observations, and has been modeled using the Sersic profile (Usher et al., 2013; Kartha et al., 2014; Hudson & Robison, 2018), the modified Hubble profile (Caso et al., 2019; De Bórtoli et al., 2022; Caso et al., 2024), and the power-law profile (Hudson & Robison, 2018). The profile shows some flattening in the inner region, which is due to the tidal disruption of the in-situ component in the slow halo assembly phase (see §3.6). The disruption changes the profile of the GC distribution, reduces its amplitude and makes its shape less concentrated, in comparison with the profile of all GCs selected at their formation (black dashed curve). As disrupted GCs can contribute to the diffuse stellar component, a part of the metal-poor and alpha-enhanced stellar population observed in the Milky Way may have originated from the disrupted GCs (Wallerstein, 1962; Weinberg et al., 2019; Rojas-Arriagada et al., 2019; Lian et al., 2020). Separated into Pop-I and Pop-II populations, the profile of Pop-I (red curve) is more centrally concentrated than that of the Pop-II (blue curve), which owes to the extension of the profile caused by metal-poor GCs brought in by mergers. This prediction is consistent with the observed negative metallicity gradient of GCs in most galaxies (e.g. Usher et al., 2013; Fahrion et al., 2020a; Harikane et al., 2023). The outer profiles ($r\sim 100\,{\rm{kpc}}$) for both Pop-I and Pop-II GCs in massive galaxies (right panel) are very similar, indicating a flat metallicity gradient in the outskirt of massive galaxies. This is indeed seen in the BCG sample of Harris (2023, see their figure 24).
@@ 4.8 Globular clusters at high redshift
The observation of GCs usually requires high spatial resolution so that GCs can be separated from other point(-like) sources. This is typically feasible only at low redshift, and becomes challenging at high redshift as the angular sizes involved are small. However, under certain conditions, GCs can be observed at high redshift. One promising way is to rely on gravitational strong lensing, which can magnify and stretch background images, making small sources observable and spatially distinguishable. Such observations have indeed carried out for a number of young and massive star clusters with sizes comparable to those of GCs up to $z\approx 10$ (e.g. Vanzella et al., 2022, 2023; Welch et al., 2023; Lin et al., 2023; Claeyssens et al., 2023; Adamo et al., 2024; Messa et al., 2024; Fujimoto et al., 2024; Mowla et al., 2024). The detectability of GCs can probably be enhanced further if the initial mass function (IMF) relevant for high-$z$ GCs is top-heavy, as suggested by some theoretical studies (Larson, 1998; Raiter et al., 2010; Marks et al., 2012; Haghi et al., 2020; Chon et al., 2021; Sharda & Krumholz, 2022) and hinted by some analyses based on observations (Zaritsky et al., 2014; Zhang et al., 2018; Dib et al., 2022; Upadhyaya et al., 2024). Here we make some predictions for the GC population at high redshift, and provide some guidance for future observations.
@@ 4.8.1 Connection to host galaxies and halos
The second column of Fig. 16 shows the relation of the total mass of GCs to the stellar mass of the host galaxy (GC + non-GC). Interestingly, the total mass of all GCs, $M_{*}^{\text{(GCs)}}$ shown by black curves, is proportional to $M_{*}$ at $z\gtrsim 4$. At lower $z$, the decrease of $M_{*}^{\text{(GCs)}}/M_{*}$ with decreasing redshift appears to be the fastest at around the Milky Way mass, producing a non-linear relation between the two masses. A separation of $M_{*}^{\text{(GCs)}}$ into contributions from Pop-I (red) and Pop-II (blue), plotted in the lower middle panel, shows that the linear $M_{*}^{\text{(GCs)}}-M_{*}$ relation at high $z$ seen in the upper middle panel originates from the balance between the two formation channels in halos of different masses (see Fig. 7). A more quantitative analysis can be made using $n_{\rm sgc}\mathcal{M}_{2}^{2}$ given by Eq. (3.1). As shown there, the high-density tail of the sub-cloud distribution depends only weakly on halo mass (via $f_{\rm gas}$ and $f_{\rm str}$) in the SN-driven regime, and so the fraction of stellar mass in GCs depends strongly only on the redshift. At lower redshift, however, two channels diverge, with the Milky Way mass falling in the gap between the two channels (shown by the blank region of Fig. 7). For comparison, we show the results of Bastian et al. (2020) as symbols in the upper central panel. Quite different to our results, the linear relation between $M_{*}^{\text{(GCs)}}$ and $M_{*}$ in their results extends to $z=0$. They argued that this linear relation is more physical than that between $M_{*}^{\text{(GCs)}}$ and $M_{\rm v}$. Since the largest difference between the predictions is at $z\sim 0$, the two models can be distinguish by measuring $M_{*}^{\text{(GCs)}}$ in galaxies with $M_{*}$ between $10^{9}$ and $10^{11}{\rm M}_{\odot}$. The third column of Fig. 16 is similar to the second column, but with $M_{*}$ replaced by $M_{\rm*,bulge}$, the bulge mass of galaxies. Since the bulge of a galaxy typically represents the dynamically hot stellar component of the galaxy, its mass may be more closely related to $M_{*}^{\text{(GCs)}}$, as both are affected by early violent collapse of the halo and major mergers experienced by the host galaxy. This is indeed seen in the right column, where the mass ratios at $z\lesssim 2$ become much flatter for galaxies with $M_{*}>10^{10}{\rm M}_{\odot}$ and for the total, Pop-I and Pop-II populations.
@@ 4.8.2 The cosmic formation history of globular clusters
The large fraction of stars that can form in GCs in a high-redshift galaxy has important implications. Indeed, if an observed high-$z$ galaxy is not properly resolved and its light is dominated by young GCs, the properties of the galaxy may not be interpreted correctly in terms of the picture we have from local galaxies. For example, the Sersic index, which is traditionally used as a measure of the concentration of light in a galaxy with a smooth profile, does not have a clear meaning for a galaxy made up of a set of star clusters. A basic challenge in future observations of high-$z$ galaxies is, therefore, to resolve galaxies down to the scales of GCs so as to identify star clusters reliably in a given field. To this end, our model, which predicts the stellar mass, star formation rate, and size of individual GCs at different redshifts in different halos, can be used to predict observable quantities to explore in detail the feasibility of observing high-$z$ GCs in current and future observations. We will come back to this in a future paper.
@@ 5 Summary and discussion
    There are two channels to make a sub-cloud sufficiently dense to form a GC. The first is shock compression in a turbulent medium produced by gravity and/or SN feedback. This can deform the log-normal density distribution of sub-clouds, producing a high-density, power-law tail, $n_{\rm sc}\gtrsim n_{\rm sgc}\mathcal{M}^{2}$, determined by the Mach number ${\cal M}$3.1). The second channel is through radiative cooling in metal-poor sub-clouds ($Z\lesssim 0.02Z_{\odot}$), where cloud fragmentation and star formation are prohibited before a sub-cloud reaches a high density $n_{\rm sf,1}$3.2). Both channels are able to lift the density of a sub-cloud to the SN-free regime, $n_{\rm sc}^{\prime}\geqslant n_{\rm snf}\approx 10^{3.5}\,{\rm cm}^{-3}$, to avoid disruption by SN feedback before forming a GC.
@@
    A number of predictions are made for GCs at high redshift. These include the relations between $M_{*}^{\text{(GCs)}}$ and the host mass (host halo mass, total stellar mass and bulge mass of the host galaxy), and the cosmic formation history of GCs in halos of different final masses.
@@ Acknowledgements
YC is funded by the China Postdoctoral Science Foundation (grant No. 2022TQ0329). This work is also supported by the National Natural Science Foundation of China (NSFC, Nos. 12192224, 11733004 and 11890693) and CAS Project for Young Scientists in Basic Research (grant No. YSBR-062). YC thanks Kai Wang, Fangzhou Jiang, Tao Wang, Yu Rong and Enci Wang for their valuable insights and discussions, and thanks Yi Mao and Chen Chen for their technical support. The authors would like to express their gratitude to the Tsinghua Astrophysics High-Performance Computing platform at Tsinghua University and the Supercomputer Center of the University of Science and Technology of China for providing the necessary computational and data storage resources that have significantly contributed to the research results presented in this paper.
@@
Credit for the images of the structures at the top of Fig. 1: Herschel image of gas sub-clouds around young stars by ESA/NASA/JPL-Caltech/AURA/NSF/STScl/Univ. of Toledo; HST and JWST images of the NGC 628 galaxy by NASA/ESA/CSA/STScI/Janice Lee (STScI)/Thomas Williams (Oxford); Simulated image of halos and cosmic large-scale structures by Li et al. (2022).
@@
The code repository TwoPhaseGalaxyModel [^5] https://github.com/ChenYangyao/two-phase-galaxy-model implements the model described in this series of papers. All data used in this paper, including data points displayed in figures and observational results used for calibration and comparison, will be distributed along with the repository. The GC catalog of MW and M31 produced by the empirical model of Chen & Gnedin (2024) is available online [^6] https://github.com/ognedin/gc_model_mw. The data of the IllustrisTNG project can be found at their website [^7] https://www.tng-project.org/.

### v2

@@ 4.6 Spatial distribution
Figure 15: Projected spatial distribution of GCs within a MW-size galaxy with $M_{\rm v}=10^{12}\,h^{-1}{\rm M_{\odot}}$ at $z=0$. All galactic GCs predicted by our model and resampled with the BSampling method (see §4.1) are shown. Each dot represents a survived GC, with bold and thin edges indicating their in-situ and ex-situ origins, respectively. The size of each circle is proportional to $R_{\rm cls}$, the half-mass radius of the GC (see §3.4). Metallicity of each GC is color-coded according to the color bar on the top. The big circle indicates two times the exponential truncation radius of the in-situ, slow-phase GC system of this galaxy. See §4.6 for a detailed discussion of this figure.
Figure 16: Upper row: number density profile of galactic GCs. Lower row: surface number density profile of galactic GCs. Three columns show the results for central galaxies with different halo masses at $z=0$. Both in-situ and ex-situ components are included. In each panel, solid black, red and blue curves are for all, Pop-I and Pop-II GCs, respectively, all including only survived GCs. Dashed black curve includes all GCs at formation (see §4.1 for the selection). Each curve shows the median of GC systems, with error bars or shaded areas indicating the 1-$\sigma$ range. Green curves are obtained by Rodriguez et al. (2023) using the model of GC formation developed by Grudić et al. (2023) and a star-by-star model of cluster evolution, applied to a MW-size galaxy in the FIRE-2 simulation. The solid and dashed curves include their GCs at birth and survived at $z=0$, respectively. Orange markers are obtained by their compilation of MW GCs based on the catalog of Harris (1996, 2010 edition). Pink markers are staked results for 7 sets of dwarf galaxies, taken from the observations by De Bórtoli et al. (2022, including those obtained earlier by ). Purple markers are stacked results for 4 BCGs, taken from the observations by Dornan & Harris (2023). See §4.6 for a detailed discussion of this figure.
@@
where GCs belonging to the same component share a common parameter $r_{\rm t}$ that describes the characteristic size of the distribution of the GC component. The isothermal form, $1/r^{2}$, is motivated by the expectation that the fragmentation of an SGC is associated with effective cooling that keeps the temperature to be $10^{4}\,{\rm K}$ at the formation of dense sub-clouds (§3.1). The exponential truncation ensures that the distribution of galactic GCs has a finite extent. Once $r$ is sampled for a GC, the angular location is randomly drawn from a uniform distribution on a unit sphere.
@@
Upper panels of Fig. 16 shows the number density profiles of GCs in central galaxies with different halo masses at $z=0$. For the profile of all survived GCs (black solid curve), the amplitude increases rapidly with halo mass, as a result of the tight, linear relation between $M_{*}^{\rm(GCs)}$ and $M_{\rm v}$ described in §4.4. In contrast, the shape of the profile is independent of halo mass, which is a consequence of our model to assign positions to GCs (see Eq. 74). Such a universal profile has been found in observations, and has been modeled using the Sérsic profile (Usher et al., 2013; Kartha et al., 2014; Hudson & Robison, 2018), the modified Hubble profile (Caso et al., 2019; De Bórtoli et al., 2022; Caso et al., 2024), and the power-law profile (Hudson & Robison, 2018). The profile shows some flattening in the inner region, which is due to the tidal disruption of the in-situ component in the slow halo assembly phase (see §3.6). The disruption changes the profile of the GC distribution, reduces its amplitude and makes its shape less concentrated, in comparison with the profile of all GCs selected at their formation (black dashed curve). As disrupted GCs can contribute to the diffuse stellar component, a part of the metal-poor and alpha-enhanced stellar population observed in the Milky Way may have originated from the disrupted GCs (Wallerstein, 1962; Weinberg et al., 2019; Rojas-Arriagada et al., 2019; Lian et al., 2020). Separated into Pop-I and Pop-II populations, the profile of Pop-I (red curve) is more centrally concentrated than that of the Pop-II (blue curve), which owes to the extension of the profile caused by metal-poor GCs brought in by mergers. This prediction is consistent with the observed negative metallicity gradient of GCs in most galaxies (e.g. Usher et al., 2013; Fahrion et al., 2020b; Harikane et al., 2023). The outer profiles ($r\sim 100\,{\rm{kpc}}$) for both Pop-I and Pop-II GCs in massive galaxies (right panel) are very similar, indicating a flat metallicity gradient in the outskirt of massive galaxies. This is indeed seen in the BCG sample of Harris (2023, see their figure 24).
@@
The square markers in the lower panels of Fig. 16 show the surface number density profiles of GCs in dwarf galaxies and BCGs, obtained by staking the observations of De Bórtoli et al. (2022, including those obtained earlier by ) and Dornan & Harris (2023), respectively, in comparison with our model prediction in the same ranges of halo mass. The error bars here are obtained by randomly sampling the profiles, taking into account uncertainties of their fitting parameters. For BCGs, we follow the Appendix of Dornan & Harris (2023) to correct the number of GCs fainter than the detection limit by assuming a Gaussian distribution for the magnitudes of GCs. The inner profile of observed dwarfs is flatter at $r\lesssim 1\,{\rm{kpc}}$ than the model. Part of such discrepancy may come from the uncertainty in the correction of incomplete detection, while the rest, as suggested by Caso et al. (2019), can be physically caused by the efficient tidal disruption of GCs in inner regions of galaxies (see also Bílek et al., 2019, 2024, for an interpretation with alternative gravity model). The observed GCs around BCGs have an outer profile $\Sigma\sim r^{-1}$, much shallower than the model prediction. Such a profile indicates that the integration of the GC number over radius is divergent, so that a significant contribution can come from GCs in the intracluster space due to projection. In both cases, it is unknown whether or not the simple assumption of a truncated isothermal profile in our model (Eq. 74) can capture the variation with halo mass. A forward modeling with precise recipes for GC orbits and disruption (e.g. Vitral & Boldrini, 2022; Ferrone et al., 2023) within realistic galactic environments is needed to make accurate model-observation comparison. This can be achieved by embedding GC-formation and evolution models in cosmological/zoom-in simulations (e.g. Rodriguez et al., 2023), or by parametric methods with calibrations using simulations and observations (e.g. Liang et al., 2024b).
@@ 4.8 Globular clusters at high redshift
Figure 19: Fraction of in-situ star formation rate contributed by GCs, color-coded according to the color bar, as a function of galaxy stellar mass and specific star formation rate (sSFR). Four panels include galaxies within different ranges of redshift, and each dot represents a galaxy. See §4.8 for details.
The observation of GCs usually requires high spatial resolution so that GCs can be separated from other point(-like) sources. This is typically feasible only at low redshift, and becomes challenging at high redshift as the angular sizes involved are small. However, under certain conditions, GCs can be observed at high redshift. One promising way is to rely on gravitational strong lensing, which can magnify and stretch background images, making small sources observable and spatially distinguishable. Such observations have indeed carried out for a number of young and massive star clusters with sizes comparable to those of GCs up to $z\approx 10$ (e.g. Vanzella et al., 2022, 2023; Welch et al., 2023; Lin et al., 2023; Claeyssens et al., 2023; Adamo et al., 2024; Messa et al., 2024; Fujimoto et al., 2024; Mowla et al., 2024; Whitaker et al., 2025). The detectability of GCs can probably be enhanced further if the initial mass function (IMF) relevant for high-$z$ GCs is top-heavy, as suggested by some theoretical studies (Larson, 1998; Raiter et al., 2010; Marks et al., 2012; Haghi et al., 2020; Chon et al., 2021; Sharda & Krumholz, 2022) and hinted by some analyses based on observations (Zaritsky et al., 2014; Zhang et al., 2018; Tang et al., 2021; Dib et al., 2022; Upadhyaya et al., 2024). Here we make some predictions for the GC population at high redshift, and provide some guidance for future observations.
@@ 4.8.1 Connection to host galaxies and halos
The second column of Fig. 17 shows the relation of the total mass of GCs to the stellar mass of the host galaxy (GC + non-GC). Interestingly, the total mass of all GCs, $M_{*}^{\text{(GCs)}}$ shown by black curves, is proportional to $M_{*}$ at $z\gtrsim 4$. At lower $z$, the decrease of $M_{*}^{\text{(GCs)}}/M_{*}$ with decreasing redshift appears to be the fastest at around the Milky Way mass, producing a non-linear relation between the two masses. A separation of $M_{*}^{\text{(GCs)}}$ into contributions from Pop-I (red) and Pop-II (blue), plotted in the lower middle panel, shows that the linear $M_{*}^{\text{(GCs)}}-M_{*}$ relation at high $z$ seen in the upper middle panel originates from the balance between the two formation channels in halos of different masses (see Fig. 8). A more quantitative analysis can be made using $n_{\rm sgc}\mathcal{M}_{2}^{2}$ given by Eq. (3.1). As shown there, the high-density tail of the sub-cloud distribution depends only weakly on halo mass (via $f_{\rm gas}$ and $f_{\rm str}$) in the SN-driven regime, and so the fraction of stellar mass in GCs depends strongly only on the redshift. At lower redshift, however, two channels diverge, with the Milky Way mass falling in the gap between the two channels (shown by the blank region of Fig. 8). For comparison, we show the results of Bastian et al. (2020) as symbols in the upper-center panel. Quite different to our results, the linear relation between $M_{*}^{\text{(GCs)}}$ and $M_{*}$ in their results extends to $z=0$. They argued that this linear relation is more physical than that between $M_{*}^{\text{(GCs)}}$ and $M_{\rm v}$. Since the largest difference between the predictions is at $z\sim 0$, the two models can be distinguish by measuring $M_{*}^{\text{(GCs)}}$ in galaxies with $M_{*}$ between $10^{9}$ and $10^{11}{\rm M}_{\odot}$. The third column of Fig. 17 is similar to the second column, but with $M_{*}$ replaced by $M_{\rm*,bulge}$, the bulge mass of galaxies. Since the bulge of a galaxy typically represents the dynamically hot stellar component of the galaxy, its mass may be more closely related to $M_{*}^{\text{(GCs)}}$, as both are affected by early violent collapse of the halo and major mergers experienced by the host galaxy. This is indeed seen in the right column, where the mass ratios at $z\lesssim 2$ become much flatter for galaxies with $M_{*}>10^{10}{\rm M}_{\odot}$ and for the total, Pop-I and Pop-II populations.
@@ 4.8.2 The cosmic formation history of globular clusters
In Fig. 19, we show the fraction of in-situ star formation rate contributed by GCs for individual galaxies, to highlight the variation of GC-formation efficiency in different galactic environments. In addition to the dependencies on redshift and stellar mass, star-forming environment, as characterized by specific star formation rate (sSFR, defined as SFR per unit stellar mass), also affects the fraction of stars formed in GCs. At all redshifts and stellar masses, a higher sSFR leads to a higher fraction of stars formed in GCs. This is a direct outcome of our external feedback model in the destruction of sub-clouds (§3.3), where low-mass sub-clouds are preferentially destroyed by the wind feedback and a higher SFR raises the threshold mass for sub-clouds to survive. A similar situation has been found observationally that the formation efficiency of bound star clusters appears to be higher in galaxies with higher SFR surface density, albeit with ongoing debates on the potential effects of, e.g. cluster selection criteria and sample homogeneity (see §7 of Adamo et al., 2020, for a review). Note that we have not defined bound star clusters in our model, but rather adopted a conservative selection for GCs. Meanwhile, we have not included the small-scale interactions among nearby sub-clouds via different types of feedbacks and tidal forces (e.g. Deng et al., 2024). A more complete model thus requires modeling correlations of sub-clouds in position and velocity spaces, which may be achieved by calibrations using zoom-in simulations with sufficient resolution to follow ISM physics.
The large fraction of stars that can form in GCs in a high-redshift galaxy has important implications. Indeed, if an observed high-$z$ galaxy is not properly resolved and its light is dominated by young GCs, the properties of the galaxy may not be interpreted correctly in terms of the picture we have from local galaxies. For example, the Sérsic index, which is traditionally used as a measure of the concentration of light in a galaxy with a smooth profile, does not have a clear meaning for a galaxy made up of a set of star clusters. The dependence of GC formation efficiency on the star-forming environment further complicates the observations, as intense star formation usually produces dusty ISM that can obscure the radiation from young GCs. A basic challenge in future observations of high-$z$ galaxies is, therefore, to resolve galaxies down to the scales of GCs so as to identify star clusters reliably in a given field. To this end, our model, which predicts the stellar mass, star formation rate, and size of individual GCs at different redshifts in different halos, can be used to predict observable quantities to explore in detail the feasibility of observing high-$z$ GCs in current and future observations. We will come back to this in a future paper.
@@ 5 Summary and discussion
    There are two channels to make a sub-cloud sufficiently dense to form a GC. The first is shock compression in a turbulent medium produced by gravity and/or SN feedback. This can deform the log-normal density distribution of sub-clouds, producing a high-density, power-law tail, $n_{\rm sc}\gtrsim n_{\rm sgc}\mathcal{M}^{2}$, determined by the Mach number ${\cal M}$3.1). The second channel is through radiative cooling in metal-poor sub-clouds ($Z\lesssim 0.02\,{\rm Z}_{\odot}$), where cloud fragmentation and star formation are prohibited before a sub-cloud reaches a high density $n_{\rm sf,1}$3.2). Both channels are able to lift the density of a sub-cloud to the SN-free regime, $n_{\rm sc}^{\prime}\geqslant n_{\rm snf}\approx 10^{3.5}\,{\rm cm}^{-3}$, to avoid disruption by SN feedback before forming a GC.
@@
    A number of predictions are made for GCs at high redshift. These include the relations between $M_{*}^{\text{(GCs)}}$ and the host mass (host halo mass, total stellar mass and bulge mass of the host galaxy), the cosmic formation history of GCs in halos of different final masses, and the variation of the GC formation efficiency with star-forming environment (§4.8).
@@ Acknowledgements
We thank the anonymous referees for the useful reports that significantly improve this paper. YC is funded by the China Postdoctoral Science Foundation (grant No. 2022TQ0329). This work is also supported by the National Natural Science Foundation of China (NSFC, Nos. 12192224, 11733004 and 11890693) and CAS Project for Young Scientists in Basic Research (grant No. YSBR-062). YC thanks Kai Wang, Fangzhou Jiang, Tao Wang, Yu Rong, Enci Wang and Baitian Tang for their valuable insights and discussions, and thanks Yi Mao and Chen Chen for their technical support. The authors would like to express their gratitude to the Tsinghua Astrophysics High-Performance Computing platform at Tsinghua University and the Supercomputer Center of the University of Science and Technology of China for providing the necessary computational and data storage resources that have significantly contributed to the research results presented in this paper.
@@
Credit for the images of the structures at the top of Fig. 1: Herschel image of gas sub-clouds around young stars by ESA, NASA, JPL-Caltech, AURA, NSF, STScl and Univ. of Toledo; HST and JWST images of the NGC 628 galaxy by NASA, ESA, CSA, STScI, Janice Lee (STScI) and Thomas Williams (Oxford); Simulated image of halos and cosmic large-scale structures by Li et al. (2022).
@@
The code repository TwoPhaseGalaxyModel [^5] https://github.com/ChenYangyao/two-phase-galaxy-model implements the model described in this series of papers. All data used in this paper, including data points displayed in figures and observational results used for calibration and comparison, will be distributed along with the repository. The GC catalog of MW and M31 produced by the empirical model of Chen & Gnedin (2024) is available online [^6] https://github.com/ognedin/gc_model_mw. The star cluster catalogs of LEGUS survey are available online [^7] https://archive.stsci.edu/prepds/legus/. The data of the IllustrisTNG project can be found at their website [^8] https://www.tng-project.org/.